Circular No. 8942

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 J3 (McNAUGHT) R. H. McNaught reports his discovery of another comet on CCD im-

ages obtained with the 0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation tabulated below), the object described as having a condensed coma of diameter 6′′–8′′ and a faint fan tail 30′′–40′′ long spanning 20^{o} toward p.a. 240^{o}. His images on May 11.66 UT show a coma with FWHM = 6′′ and a narrow tail 45′′ long in p.a. 245^{o}; stacked images taken by McNaught on May 12.64 show a 40′′ tail in p.a. 245^{o} that appears as a fan spanning p.a. 230^{o}–260^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, A. C. Gilmore and P. M. Kilmartin (Mt. John 1.0-m f/7.7 reflector) report that their CCD images from May 11.62–11.63 show a coma diameter of ∼ 7′′ and a narrow tail 25′′ long in p.a. 250^{o}; their exposures from May 12.71–12.72 show a tail 25′′ long in p.a. 250^{o}, a coma diameter of ∼ 10′′ with total magnitude 18.7–19.0 measured in a 12′′- diameter aperture, and a starlike condensation of mag 19.9 as measured in a 5′′-diameter aperture. N. Teamo, Punaauia, French Polynesia, finds a tail 15′′ long in p.a. 250^{o} on his images taken on May 12.55–12.58 with a 0.41-m f/8.0 reflector; J. C. Pelle measures a tail 7′ long in p.a. 240^{o} from additional images taken by Teamo on May 12.61–12.63.

2008 UT  α_{2000}  δ_{2000}  Mag.
May 10.69282  21^{h}45^{m}44.^{s}98  −42^{o}44′30′′.3  18.4

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-J64; this comet may be of short period.

T = 2008 Apr. 3.968 TT  ω = 259.^{o}341
Ω = 28.660 }2000.0
q = 3.53197 AU  i = 25.427

COMET C/2008 J2 (BESHORE) Further to IAUC 8941, J. McGaha revises his description from May 7.3

UT: 10-s exposures show a dense condensation of diameter 5′′ surrounded by a coma of diameter 11′′; fifteen stacked 60-s exposures show a sharp disk of diameter 21′′ with no diffuseness and a bright elongation extending ∼ 15′′.5 from the object’s center. Sixty stacked 10-s exposures from May 10.35 to 10.40 show a uniform disk of size 12′′ surrounded by a faint coma of diameter 30′′ and a faint 68′′ tail in p.a. 261^{o}.

2008 May 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green